OpenStudy (anonymous):

Use this to find the equation of the tangent line to the curve y=5−3x^3 at the point (3−76) and write your answer in the form: y=mx+b, where m is the slope and b is the y-intercept.

OpenStudy (anonymous):

could you help me @satellite73

OpenStudy (anonymous):

so, a derivative is only an equation to give you the tangent (aka. slope) of a a line at a particular point. In your y=mx+b equation, we need exactly such a slope... so we find the derivative of your equation: \[\frac{d}{dx} (5-3x^{3}) = -9x^{2}\] okay, so in order to determine the slope of the equation at your point (3, -76), you just need the x-coordinate! Okay, so we have a slope m in our equation y=mx+b. What do we need to determine b? Well, we need a value for x and a value for y... aren't those two value part of a coordinate? I think we have what we need to determine the value for b, and therefore determine the equation for the tangent line! : )

OpenStudy (anonymous):

\[m = -9(3)^{2}\] \[m = -81\] therefore \[y=-81x+b\] sub in our point: \[-76=-81(3) + b\] \[b=167\] therefore: \[y=-81x+167\]

OpenStudy (anonymous):

THANKS!!!